OpenStudy (anonymous):

For 2x – 3y = –10, when x = 4, What is the value of y?

OpenStudy (anonymous):

3y=2x+10 for x=4 3y=8+10 y=8/3+10

OpenStudy (anonymous):

y=18/3 y=6

OpenStudy (anonymous):

Yep, that's right. My mistake :)